An object moves along the x axis according to the equation π‘₯ = 3.00𝑑

2 βˆ’ 2.00𝑑 + 3.00,

where x is in meters and t is in seconds. Determine (a) the average speed between t = 2.00 s

and t = 3.00 s, (b) the instantaneous speed at t = 2.00 s and at t = 3.00 s, (c) the average

acceleration between t = 2.00 s and t = 3.00 s, and (d) the instantaneous acceleration at t =

2.00 s and t = 3.00 s. (e) At what time is the object at rest?


1
Expert's answer
2021-03-08T08:27:46-0500

(a) Let's first find the position of the object at "t=2\ s" and "t=3\ s":


"x_1(t=2\ s)=3.0\cdot(2)^2-2.0\cdot2+3.0=11\ m,""x_2(t=3\ s)=3.0\cdot(3)^2-2.0\cdot3+3.0=24\ m."

By the definition of the average speed, we have:


"v_{av}=\dfrac{x_2-x_1}{t_2-t_1}=\dfrac{24\ m-11\ m}{3\ s-2\ s}=13\ \dfrac{m}{s}."

(b) The speed is the derivative of the position of the object with respect to time:


"v=\dfrac{dx}{dt}=\dfrac{d}{dt}(3t^2-2t+3)=6t-2."

Let's find the instantaneous speed at "t=2\ s":


"v_1(t=2\ s)=6\cdot2-2=10\ \dfrac{m}{s}."

Let's find the instantaneous speed at "t=3\ s":


"v_2(t=3\ s)=6\cdot3-2=16\ \dfrac{m}{s}."

(c) By the definition of the average acceleration, we have:


"a_{avg}=\dfrac{v_2-v_1}{t_2-t_1}=\dfrac{16\ \dfrac{m}{s}-10\ \dfrac{m}{s}}{3\ s-2\ s}=6\ \dfrac{m}{s^2}."

(d) The acceleration is the derivative of the speed of the object with respect to time:


"a=\dfrac{dv}{dt}=\dfrac{d}{dt}(6t-2)=6\ \dfrac{m}{s^2}."

As we can see from the calculations, at "t=2\ s" and at "t=3\ s" the acceleration remains constant and equals "6\ \dfrac{m}{s^2}".

(e) The object is at rest when "v=0" :


"6t-2=0,""t=\dfrac{2}{6}=0.33\ s."